write an equation of the line that passes through (-2,3) and is perpendicular to the line y=2x+1
Answer by Mathtut(3670)

parallel lines have equal slopes but perpendicular lines are negative inverses of each other. So we need to first write the equation in y=mx+b format and we see it already is written in that format. We see from this that the slope is 2 therefore if the other line is perpendicular to this is has a negative inverse slope which would equal to -1/2. since we have the slope and a given point(-2,3) on the perpendicular line we can use the point slope formula which is y-k=m(x-h) where h,k is a point on the line. so using our given point and slope we have y-3=-1/2(x-(-2)) simplify y=-1/2x+5
answer: y = -1/2x + 5